# Boosting Open-Domain Continual Learning via Leveraging Intra-domain Category-aware Prototype

### v1

@@ Introduction
In addition to the general prompt, we provide instance-level domain prior prompt for the image encoder. Using the prototype $P_{n}$ as input, we propose an Image Cross-Attention (ICA) module to dynamically generate instance-level prompt. The class token ${cls}^{(h)}$ from the $h$th layer of the image encoder serves as the $Q$, while the domain prototype serves as $K$ and $V$, resulting in the domain prior prompt $e_{p}$, $e_{p}\in{\mathbb{R}^{1\times d^{\mathcal{I}}}}$.
$Q^{\prime}=W_{Q}^{\prime}\cdot{cls}^{(h)},K^{\prime}=W_{K}^{\prime}\cdot P_{n},V^{\prime}=W_{V}^{\prime}\cdot P_{n},$ (9)
where $W_{Q}^{\prime}\in\mathbb{R}^{d^{\mathcal{I}}\times d^{\mathcal{I}}}$, $W_{K}^{\prime}$ and $W_{V}^{\prime}\in\mathbb{R}^{d^{\mathcal{T}}\times d^{\mathcal{I}}}$ are learnable projection matrices in the image branch.
@@
Inference Stage
For test images from previously seen domains in the ODCL-CIL setting, we first employ the prototype Task-ID discriminator to determine the Task-ID of the test image, as shown in Figure 2(c). Specifically, for a test image $x$, we use the original CLIP image encoder to extract the corresponding image feature $I$.
$I=\mathcal{I}(x).$ (10)
We then calculate the cosine similarity between the image feature $I$ and the learned domain prototypes of different domains. For each domain, we select the highest similarity score, and then compare the similarities scores across different domains to obtain the Task-ID. Subsequently, we select the corresponding domain component from the Domain prior prompt pool. Classification of the test image is then performed within the category set associated with the identified Task-ID domain and domain component.
For test images from unseen domains, we adopt an approach similar to ColeCLIP. We utilize CLIP to perform zero-shot classification on these images from unseen domains. However, if the unseen dataset contains some categories that have been previously seen, we will query the previously seen domains in which the corresponding categories appear. We will then utilize the components associated with those domains to compute the cosine similarity for the respective categories. The category with the highest similarity will be selected as the corresponding result. Finally, we will compare this with the cosine similarity of CLIP for unseen categories, choosing the largest as the final classification result.
Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Mean
Last CODA-Prompt 12.81 65.38 63.49 53.83 56.91 56.64 83.31 83.99 76.83 53.65 75.13 62.00
@@
Table 1: $Last$, $Forgetting$, and $Avg$ results(%) in ODCL-CIL. CODA-Prompt and LAE are not applicable to the Avg metric.
Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Mean
Last CODA-Prompt 43.59 92.97 74.47 75.21 92.57 93.07 90.25 98.61 92.75 83.80 78.75 83.28
@@
Frogetting CODA-Prompt 47.49 92.84 78.23 76.40 94.51 95.23 90.20 98.55 93.12 83.99 78.75 84.48
@@
Table 2: $Last$, $Forgetting$, and $Avg$ results(%) in ODCL-TIL. CODA-Prompt and LAE are not applicable to the Avg metric.
Experiments
Experiments Setup
Datasets
The ODCL task comprises 11 datasets across various domains, including Aircraft (maji2013fine), Caltech101 (fei2004learning), CIFAR100 (2009Learning), DTD (cimpoi2014describing), EuroSAT (helber2019eurosat), Flowers (nilsback2008automated), Food (bossard2014food), MNIST (deng2012mnist), OxfordPet (parkhi2012cats), StanfordCars (krause20133d), and SUN397 (xiao2010sun). In the ODCL task, each dataset represents a distinct domain that needs to be addressed. The sequence in the ODCL task typically follows two orders, as previously adopted in the works of ZSCL (zheng2023preventing), ColeCLIP (li2024coleclip). The Order-I arranges the datasets alphabetically by their names, while the Order-II is determined randomly. The experimental results presented in the main text are based on the Order-I , whereas the supplementary materials provide results based on the Order-II.
Implementation Details
Building upon the previous research by ZSCL and ColeCLIP, we adopt the CLIP-ViT-B/16 (radford2021learning) model as the foundational architecture for tackling the ODCL task. All experimental settings strictly follow ColeCLIP (li2024coleclip), i.e., we employ the Adam optimizer with a fixed learning rate of $1\times 10^{-3}$, adopt a batch size of 128, and the specific number of epochs for each dataset can be found in the supplementary materials.
For the text encoder, we employ a single-layer learnable prompt as the general prompt integrated with a TSA module. The domain prior prompt generated by the TSA module replaces the original prompt in the $8$-th layer. For the image encoder, we utilize a 7-layer deep learnable prompts as general prompts, complemented by an ICA module. The instance domain prior prompt generated by the ICA module replaces the original prompt in the $8$-th layer. In line with the prompt length employed in ColeCLIP, we set the prompt length to 1 for both encoders. Similar to CLIP, the temperature $\tau$ is set to 0.01.
Evaluation Metrics
Following previous work such as ZSCL, ColeCLIP, we use Avg, Last, Transfer, and Forgetting as evaluation metrics to assess performance on the ODCL task. For the $n$th dataset $\mathcal{D}_{n}$, we use $A_{n}$, $F_{n}$ and $R_{n}$ to represent the mean accuracy, Frogetting and Transfer metrics, respectively.
Avg measures the overall classification accuracy across the 11 datasets, $Avg=\frac{1}{N}\sum_{n=1}^{N}A_{n}$, where $N$ is the total number of datasets. Specifically, after training on a given dataset, we evaluate the model’s performance on all 11 datasets (including both seen and unseen datasets) and compute the mean accuracy across these datasets, denoted as $A_{n}$, $A_{n}=\frac{1}{N}\sum_{u=1}^{N}A_{u}^{n}$, $A_{u}^{n}$ is the accuracy of the model on the $n$th dataset after training in $u$th dataset. After training on all 11 datasets, we calculate the mean of all accuracies to obtain Avg. Last refers to the final accuracy across the 11 datasets after training on the final task. Forgetting measures the degree of performance degradation on previously seen datasets. The computation involves calculating the mean accuracy for each previously seen dataset and then averaging these values to obtain the Forgetting metric, $Frogetting=\frac{1}{N}\sum_{n=1}^{N}F_{n}$ and $F_{n}=\frac{1}{N-n+1}\sum_{v=n}^{N}A_{v}^{n}$. Transfer assesses the model’s zero-shot capability on unseen datasets. It is computed by averaging the zero-shot accuracies of the datasets before the model is trained on them, $R_{n}=\frac{1}{n-1}\sum_{v=1}^{n-1}A_{v}^{n}$. The mean of these 10 values constitutes the Transfer metric, $Transfer=\frac{1}{N-1}\sum_{n=2}^{N}R_{n}$.
@@
Table 5: Ablation analysis on the text branch. CoOp represents learnable prompt in $1$th layer. LP is learnable prompt in $8$th layer. TSA represents domain prior prompt in $8$th layer.
@@
Table 6: Ablation analysis on the image branch. VPT represents learnable prompt in 1-7 layers. LP is learnable prompt in $8$th layer. ICA represents instance-level domain prior prompt in $8$th layer.
Figure 3: (a) represents the influence of different prompt length and (b) represents the influence of different prompt replacement depth.
Comparison Methods
We selects CLIP (radford2021learning), ZSCL, MoE-Adapters (yu2024boosting), and ColeCLIP as the baseline methods for comparison. ZSCL represents the pioneering approach for ODCL task, MoE-Adapters explore the application of the MoE structure in the ODCL task while ColeCLIP builds upon it by further refining designs and enhancing performance. Since the MoE-Adapters does not fully align with this task, we modified its implementation to meet the demands of the ODCL task. The specific implementation details are provided in the supplementary materials. Following the methodology adopted by ColeCLIP, we also incorporate traditional continual learning methods, specifically CODA-Prompt (smith2023coda) and LAE (gao2023unified), for comparative analysis. However, since these methods are designed for closed-domain scenarios and lack zero-shot capabilities, they are only applicable for evaluating the Last and Forgetting metrics, and not for assessing the Avg and Transfer metrics.
Main Properties
Performance on ODCL-CIL
As previously mentioned, ODCL-CIL presents a more challenging task compared to ODCL-TIL due to the absence of Task-ID information during the inference stage. Table 1 provides a comparison between our proposed method DPeCLIP and other approaches, highlighting significant improvements in performance.
Notably, DPeCLIP achieves improvements of 4.90% and 3.33% over CoLeCLIP in the Last and Forgetting metrics, respectively. For the Caltech101 and CIFAR100 datasets, which are significantly affected by the category-relatedness issue, the first category of methods, ZSCL and ColeCLIP, exhibited noticeable forgetting phenomena. MoE-Adapters and our method demonstrate that a robust Task-ID discriminator can effectively address the category-relatedness problem. Additionally, due to the shortcomings of MoE-Adapters’ Task-ID discriminator, the performance on many datasets is significantly affected, such as Caltech101, EuroSAT, and Food.
Our method shows an enhancement of 2.37% compared to CoLeCLIP in Avg. Although MoE-Adapters demonstrates strong performance on the DTD task, it utilizes the AutoEncoder as the Task-ID discriminator, which fails to recognize test images from EuroSAT. Thus, it negatively impacts the overall performance. In comparison, our prototype Task-ID discriminator approach demonstrates exceptional performance in terms of Task-ID identification accuracy, which ultimately enhances the overall model performance. Overall, our method demonstrates robust performance, consistently outperforming previous approaches across the majority of datasets.
@@
Table 7: The impact of prototype granularity for the Task-ID discriminator. Domain-level granularity represents one prototype for each domain, while category-level granularity represents one prototype for each category within each domain.
@@
Table 8: The impact of different types of prototypes on the Task-ID discriminator. TP represents text prototypes, while IP represents image prototypes.
@@
Table 9: The impact of different type of prototype for domian prior prompt. TP is referring text prototype and IP is referring image prototype.
Figure 4: Comparison of t-SNE between MoE-Adapters and our method for the Task-ID discriminator’s output.
Performance on ODCL-TIL
For the ODCL-TIL task, DPeCLIP continues to demonstrate exceptional performance, as illustrated in Table 2. In terms of the Last, Forgetting, and Avg metrics, DPeCLIP achieves the highest results, surpassing CoLeCLIP by 2.18%, 1.74%, and 1.14%, respectively. This highlights that DPeCLIP not only excels in addressing the issue of forgetting but also achieves superior learning outcomes across various domain tasks. DPeCLIP effectively balances the trade-off between learning capability and forgetting mitigation, resulting in outstanding overall performance.
Transfer performance
Table 3 presents the Transfer metric for models on the ODCL-TIL task. ZSCL, by fully fine-tuning the CLIP model, experiences a significant decline in transfer performance, with a reduction of 1.21%. CoLeCLIP mitigates this issue by maintaining a large vocabulary, resulting in notable improvements. MoE-Adapters utilizes a mixture of expert structure to retain CLIP’s original knowledge, and it exhibits relatively good performance on the transfer metric. In contrast, DPeCLIP introduces domain prototypes as input for the domain prior prompts, aiming to retain the original knowledge of CLIP as much as possible, thereby concurrently outperforming other methods on various other metrics, achieving the same level of performance on the Transfer metric as MoE-Adapters, with only a 0.17% reduction.
Analysis and discussion
Module Analysis
We conduct comprehensive ablation experiments on DPeCLIP to demonstrate the impact of different components of our proposed DPeCLIP. Since the ODCL-CIL task can better evaluate the overall performance of the model, Table 4 shows the performance using different components on ODCL-CIL task. The first row represents the performance of the original CLIP on ODCL-CIL task. By incorporating the prototype Task-ID discriminator, without requiring any training, the Avg, Last, and Forgetting metrics show significant improvements, 3.06%, 5.75% and 4.8% respectively. This method effectively addresses the issue of category-relatedness across different domains.
TC (Text Component) refers to the general prompt and domain prior prompt generated by the TSA module for the text encoder and IC (Image Component) represents the general prompt and instance-level domain prior prompt generated by the ICA module for the image encoder. Compared to IC, the TC shows a more significant increase in the three main metrics, demonstrating that fine-tuning the text component of CLIP is more effective. Additionally, the domain prior prompt generated from the domain prototype performs well in retaining the original knowledge of CLIP, resulting in only a slight decline in the Transfer metric. The integration of the three modules achieves the best performance in Avg, Last, and Forgetting metrics, with only a very slight decline in the Transfer metric. The detailed ablation experiments for each component are shown below.
Detailed ablation study of TC
Table  5 shows a detailed analysis of TC. CoOp represents the general prompt for the first layer, LP refers to the learnable prompt for the $8$th layer, and TSA module provides the domain prior prompt for the $8$th layer. Compared to the combination of CoOp and LP, the combination of CoOp and TSA shows significant improvements across all metrics, demonstrating the effectiveness of the domain prior prompt for the text encoder.
Detailed ablation study of IC
Table 6 presents a detailed analysis of IC. VPT represents the learnable prompt for the first $7$ layers used as the general prompt, LP refers to the learnable prompt for the $8$th layer, and ICA stands for the instance-level domain prior prompt generated by the ICA module for the $8$th layer. Notably, the combination of VPT and ICA shows significant growth not only in all metrics, including the transfer metric, demonstrating that the domain prior prompt retains more of CLIP’s original knowledge compared to the learnable prompt.
Prompt Length
Figure 3(a) presents the performance of DPeCLIP with different prompt length on ODCL task. Given the same settings for hyperparameters such as training iterations and learning rate, DPeCLIP exhibits the best performance across all metrics when the prompt length is 1.
Replacement depth
Figure 3(b) presents the replace depths of the domain prior prompt obtained using the TSA and ICA modules. We conducted experiments by using layers 5 to 11 of the CLIP encoder as the replacement depth. Although the Transfer metric achieves the best performance when inserted at layer 6, a comprehensive evaluation leads us to ultimately select layer 8, as it yields the best overall performance. We hypothesize that inserting the layer 8 can enable the model to better balance the consideration of low-level and high-level features.
@@
Table 7 shows the impact of different prototype granularities on the Task-ID discriminator. Compared to domain-level prototypes, category-level prototypes can retain more domain information, resulting in more accurate Task-ID judgments and better model performance.
@@
Table 8 shows the impact of different types of prototypes on the Task-ID discriminator. Compared to the confusion issues associated with TP, IP better represents domain-level information, resulting in improved performance of the Task-ID discriminator. Additionally, we found that using IP and TP together can slightly enhance overall performance.
@@
Table 9 illustrates the influence of different type of prototype for domain prior prompt. It can be observed that using either type of prototype individually can achieve good performance. However, since combining both yields the best results, we ultimately choose to use a combination of TP and IP.
@@
Figure 4 shows the different distributions between MoE-Adapters and our method in the Task-ID discriminator’s output. MoE-Adapters exhibit significant confusion among different datasets, whereas our method effectively reduces this confusion. The comparison of the magnified sections further demonstrates the superior performance of our method on the Task-ID discriminator task.

### v2

@@ Introduction
In addition to the general prompt, we provide instance-level domain prior prompt for the image encoder. Using the prototype $P_{n}$ as input, we propose an Image Cross-Attention (ICA) module to dynamically generate instance-level prompt. The class token ${cls}^{(h)}$ from the $h$th layer of the image encoder serves as the $Q^{\prime}$, while the domain prototype serves as $K^{\prime}$ and $V^{\prime}$ through the formula $Q^{\prime}=W_{Q}^{\prime}\cdot{cls}^{(h)},K^{\prime}=W_{K}^{\prime}\cdot P_{n},V^{\prime}=W_{V}^{\prime}\cdot P_{n}$, where $W_{Q}^{\prime}\in\mathbb{R}^{d^{\mathcal{I}}\times d^{\mathcal{I}}}$, $W_{K}^{\prime}$ and $W_{V}^{\prime}\in\mathbb{R}^{d^{\mathcal{T}}\times d^{\mathcal{I}}}$ are learnable projection matrices in the image branch, yielding the domain prior prompt $e_{p}$, where $e_{p}\in{\mathbb{R}^{1\times d^{\mathcal{I}}}}$.
@@
Task Method Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Mean
ODCL-CIL Last CODA-Prompt 12.81 65.38 63.49 53.83 56.91 56.64 83.31 83.99 76.83 53.65 75.13 62.00
@@
ODCL-TIL Last CODA-Prompt 43.59 92.97 74.47 75.21 92.57 93.07 90.25 98.61 92.75 83.80 78.75 83.28
@@
Forgetting CODA-Prompt 47.49 92.84 78.23 76.40 94.51 95.23 90.20 98.55 93.12 83.99 78.75 84.48
@@
Table 1: Last, Forgetting, and Avg (%) in ODCL-CIL and ODCL-TIL. CODA-Prompt and LAE are not applicable to Avg.
Inference Stage
For a test image from previously seen domains in ODCL-CIL, we first employ the prototype Task-ID discriminator to determine its Task-ID, as shown in Figure 2(c). For a test image $x$, we use $\mathcal{I}(x)$ to extract the image feature $I$. We then calculate the cosine similarity between $I$ and the learned domain prototypes of different domains. For each domain, we select the highest similarity score, and then compare the similarities scores across different domains to obtain the Task-ID. Next, we select the corresponding domain component from the domain prior prompt pool. Classification of the test image is then performed within the category set associated with the identified Task-ID domain and domain component.
For a test image from unseen domains, we adopt an approach similar to ColeCLIP. We use CLIP to perform zero-shot classification on the image from unseen domains. However, if the unseen dataset contains categories that have been previously seen, we will query the previously seen domains in which the corresponding categories appear. We then use the components associated with those domains to compute the cosine similarity for the respective categories. The category with the highest similarity is selected as the result. Finally, we compare this with the cosine similarity of CLIP for unseen categories, choosing the largest as the final result.
@@
Experiments
Experiments Setup
Datasets
The ODCL task comprises 11 datasets across various domains, including Aircraft (Maji et al. 2013), Caltech101 (Fei-Fei, Fergus, and Perona 2004), CIFAR100 (Krizhevsky and Hinton 2009), DTD (Cimpoi et al. 2014), EuroSAT (Helber et al. 2019), Flowers (Nilsback and Zisserman 2008), Food (Bossard, Guillaumin, and Van Gool 2014), MNIST (Deng 2012), OxfordPet (Parkhi et al. 2012), StanfordCars (Krause et al. 2013), and SUN397 (Xiao et al. 2010). The sequence in the ODCL task typically follows two orders, as previously adopted in the works of ZSCL (Zheng et al. 2023), ColeCLIP (Li et al. 2024). The Order-I arranges the datasets alphabetically by their names, while the Order-II is determined randomly. The experimental results presented in the main text are based on the Order-I , whereas the supplementary materials provide results based on the Order-II.
Implementation Details
Building upon the previous research by ZSCL and ColeCLIP, we adopt the CLIP-ViT-B/16 (Radford et al. 2021) model as the foundational architecture for the ODCL task. All experimental settings strictly follow ColeCLIP (Li et al. 2024), more detailed descriptions are provided in the supplementary materials.
For the text encoder, we employ a single-layer learnable prompt as the general prompt integrated with a TSA module. The domain prior prompt generated by the TSA module replaces the original prompt in the $8$-th layer. For the image encoder, we utilize a 7-layer deep learnable prompts as general prompts, complemented by an ICA module. The instance-level domain prior prompt generated by ICA replaces the original prompt in the $8$-th layer. In line with the prompt length employed in ColeCLIP, we set the prompt length to 1 for both encoders.
Evaluation Metrics
Following previous work such as ZSCL, ColeCLIP, we use Avg, Last, Transfer, and Forgetting as evaluation metrics to assess performance on the ODCL task. For the $n$th dataset $\mathcal{D}_{n}$, we use $A_{n}$, $F_{n}$ and $R_{n}$ to represent the mean accuracy, Forgetting and Transfer metrics, respectively. More detailed descriptions are provided in the supplementary materials.
@@
Table 4: Analysis on the text branch. CoOp represents learnable prompt in $1$th layer. LP is learnable prompt in $8$th layer. TSA represents domain prior prompt in $8$th layer.
Comparison Methods
We selects CLIP (Radford et al. 2021), ZSCL, MoE-Adapters (Yu et al. 2024), CODA-Prompt(Smith et al. 2023), LAE (Gao et al. 2023) and ColeCLIP as the baseline methods for comparison. More detailed descriptions can be found in the supplementary materials.
Main Properties
Performance on ODCL-CIL
As previously mentioned, ODCL-CIL presents a more challenging task compared to ODCL-TIL due to the absence of Task-ID information during the inference stage. Table 1 ODCL-CIL part provides a comparison between our proposed method DPeCLIP and other approaches, highlighting significant improvements in performance.
Notably, DPeCLIP achieves improvements of 4.90% and 3.33% over CoLeCLIP in the Last and Forgetting metrics, respectively. For Caltech101 and CIFAR100, which are significantly affected by the category-relatedness issue, the first category of methods, ZSCL and ColeCLIP, exhibited noticeable forgetting phenomena. MoE-Adapters and our method demonstrate that a robust Task-ID discriminator can effectively address the category-relatedness problem. Additionally, due to the shortcomings of MoE-Adapters’ Task-ID discriminator, the performance on many datasets is significantly affected, such as Caltech101, EuroSAT, and Food.
Our method shows an enhancement of 2.37% compared to CoLeCLIP in Avg. Although MoE-Adapters demonstrates strong performance on the DTD task, it utilizes the AutoEncoder as the Task-ID discriminator, which fails to recognize test images from EuroSAT. Thus, it negatively impacts the overall performance. In comparison, our prototype Task-ID discriminator approach demonstrates exceptional performance in terms of Task-ID identification accuracy, which ultimately enhances the overall model performance.
@@
Table 5: Analysis on the image branch. VPT represents learnable prompt in 1-7 layers. LP is learnable prompt in $8$th layer. ICA represents domain prior prompt in $8$th layer.
@@
Table 6: The impact of prototype granularity for the Task-ID discriminator in ODCL-CIL task. Domain-level represents one prototype for each domain, while category-level represents one prototype for each category within each domain.
Performance on ODCL-TIL
For the ODCL-TIL task, DPeCLIP continues to demonstrate exceptional performance, as illustrated in Table 1 ODCL-TIL part. In terms of the Last, Forgetting, and Avg metrics, DPeCLIP achieves the highest results, surpassing CoLeCLIP by 2.18%, 1.74%, and 1.14%, respectively. This highlights that DPeCLIP not only excels in addressing the issue of forgetting but also achieves superior learning outcomes across various domain tasks. DPeCLIP effectively balances the trade-off between learning capability and forgetting mitigation.
Transfer Performance
Table 2 presents the Transfer metric for models on the ODCL-TIL task. ZSCL, by fully fine-tuning the CLIP model, experiences a significant decline in transfer performance, with a reduction of 1.21%. CoLeCLIP mitigates this issue by maintaining a large vocabulary, resulting in notable improvements. MoE-Adapters utilizes a MoE structure to retain CLIP’s original knowledge, and it exhibits relatively good performance on the transfer metric. In contrast, DPeCLIP introduces domain prototypes as input for the domain prior prompts, aiming to retain the original knowledge of CLIP as much as possible, thereby concurrently outperforming other methods on various other metrics, achieving the same level of performance on the Transfers metric as MoE-Adapters, with only a 0.17% reduction.
Analysis and Discussion
Module Analysis
We conduct comprehensive ablation experiments on DPeCLIP to assess the impact of different components. Since the ODCL-CIL task better evaluates overall model performance, Table 3 presents results using various components for this task. The first row shows the performance of the original CLIP. By incorporating the prototype Task-ID discriminator without additional training, the Avg, Last, and Forgetting metrics improve significantly by 3.06%, 5.75%, and 4.8%, respectively. This method effectively addresses category-relatedness issues across domains.
TC (Text Component) refers to the general prompt and domain prior prompt generated by the TSA module for the text encoder and IC (Image Component) represents the general prompt and instance-level domain prior prompt generated by the ICA module for the image encoder. Compared to IC, the TC shows a more significant increase in the three main metrics, demonstrating that fine-tuning the text encoder is more effective. Additionally, the domain prior prompt generated from the domain prototype performs well in retaining the original knowledge of CLIP, resulting in only a slight decline in the Transfer metric. The integration of the three modules achieves the best performance the three main metrics, with only a very slight decline in the Transfer metric. The detailed ablation experiments are presented below.
Effectiveness of TSA
Table  4 shows an analysis of TC. Compared to the combination of CoOp and LP, the combination of CoOp and TSA shows significant improvements across all metrics, demonstrating the effectiveness of the domain prior prompt for the text encoder.
Effectiveness of ICA
Table 5 presents an analysis of IC. Notably, the combination of VPT and ICA shows significant growth in all metrics , demonstrating that the domain prior prompt retains more of CLIP’s original knowledge compared to the learnable prompt.
Figure 3: (a) represents the influence of prompt length and (b) represents the influence of prompt replacement depth.
@@
Table 7: Different types of prototypes for the Task-ID discriminator. TP is text prototype and IP is image prototype.
Prompt Length
Figure 3(a) presents the performance of DPeCLIP with different prompt length. Given the same settings for hyperparameters, DPeCLIP exhibits the best performance across all metrics when the prompt length is 1.
Replacement Depth
Figure 3(b) shows the effects of replacement depth for the domain prior prompt. We experimented with layers 5 to 11 of the encoder as replacement depth. Although the Transfer metric performs best when inserted at layer 6, we ultimately chose layer 8 for the best overall performance. We hypothesize that layer 8 allows the model to better balance low-level and high-level features.
@@
Table 8: Different types of prototype for domain prior prompt. TP is text prototype and IP is image prototype.
Figure 4: Comparison of t-SNE between MoE-Adapters and our method for the Task-ID discriminator.
@@
Table 6 shows the impact of different prototype granularities for the Task-ID discriminator. Compared to domain-level prototypes, category-level prototypes can retain more domain information, resulting in more accurate Task-ID judgments and better model performance.
@@
Table 7 shows the impact of different types of prototypes on the Task-ID discriminator in ODCL-CIL task. Compared to the confusion issues associated with TP, IP better represents domain-level information, resulting in improved performance of the Task-ID discriminator. We found that using IP and TP together can slightly enhance performance, maybe due to complementarity of different prototype.
@@
Table 8 illustrates the influence of different type of prototype for domain prior prompt. It can be observed that using either type of prototype individually can achieve good performance. To provide more comprehensive information, we adopted a combination of TP and IP, achieving the best results.
@@
Figure 4 shows the different distributions between MoE-Adapters and our method for the Task-ID discriminator. MoE-Adapters exhibit significant confusion, whereas our method effectively reduces this confusion. The comparison of the magnified sections further demonstrates the superior performance of DPeCLIP on the Task-ID discriminator task.